OpenStudy (anonymous):

Simplify: (SecX^2-TanX^2)/Cos(-X) For my lovely @wet and @shelbygt520 That Latex stuff seemed to take too long, so I hope you get that.

OpenStudy (anonymous):

For this I got the answer as CosX, but the correct answer is CosX^2.

OpenStudy (anonymous):

https://www.youtube.com/watch?v=TfzIObd_7Hw

OpenStudy (anonymous):

What the heck did I just watch?

OpenStudy (anonymous):

You just watched the answer

OpenStudy (anonymous):

ok let me explain it word by word bib help me out here

OpenStudy (anonymous):

to all of your problems

OpenStudy (anonymous):

But it wasn't the same question O.o

OpenStudy (anonymous):

(secx)^2-(tanx)^2 that's definitely your question

OpenStudy (anonymous):

oh wait there's a cos(-x)

OpenStudy (anonymous):

Oh, you mean without the denominator?

OpenStudy (anonymous):

>.>

OpenStudy (anonymous):

k imma start first shelbs

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Recall the basic trig identity \[\large \sin^2x+\cos^2x = 1\]we can then derive:\[\large \sec^2x= 1+tan^2x \]\[\large \sec^2x-tan^2x= 1 \]

OpenStudy (anonymous):

Yeah, those are the Pythogorean Identities.

OpenStudy (anonymous):

pls\[\huge \frac{(\sec^2x-\tan^2x)}{\cos(-X)}\]\[\huge (\sec^2x-\tan^2x)=1\]

OpenStudy (anonymous):

Holy...

OpenStudy (anonymous):

So the Answer is 1/Cos(-X)?

OpenStudy (anonymous):

I think there's an identity for the bottom. cos(-x) = cosx

OpenStudy (anonymous):

at least that's what it looks like from this http://www.wolframalpha.com/input/?i=cos+-x

OpenStudy (anonymous):

OOh yeah i have that identity in my book, so 1/CosX?

OpenStudy (anonymous):

aka ?

OpenStudy (anonymous):

SecX?

OpenStudy (anonymous):

that's the simplest form

OpenStudy (anonymous):

Wait, so why did my teacher put a huge red X and tell me the answer is CosX^2?

OpenStudy (anonymous):

no clue but bib is correct

OpenStudy (anonymous):

I don't think I'm wrong...

OpenStudy (anonymous):

I also don't think you're wrong, but i don't understand why the teacher did that

OpenStudy (anonymous):

Weird.

OpenStudy (anonymous):

even wolfram says it evaluates to secx(sec^2x-tan^2x) = secx http://www.wolframalpha.com/input/?i=%28SecX%5E2-TanX%5E2%29%2FCos%28-X%29

OpenStudy (anonymous):

Go up to your teacher and ask why. He probably made a mistake. "Everybody makes mistakes" - hannah montana

OpenStudy (anonymous):

your teacher may have missed something take it back to her and explain how you got the answer and show her evidence to support it

OpenStudy (anonymous):

Okay, again, thank you guys sooooo much!

OpenStudy (anonymous):

me and bib here are always here to help! :) bib where are you?! a strip club?! @we

OpenStudy (anonymous):

@wet

OpenStudy (anonymous):

Okay guys, sorry, I know I'm asking too much, but I just have this last confusing question

OpenStudy (anonymous):

woah I don't do that yo. don't sully my name @shelbygt520 bby

OpenStudy (anonymous):

you know the drill. new question tag me/her/us

OpenStudy (anonymous):

lol